What is the mole fraction of 25% glucose solution? Find its molal concentration
1
Expert's answer
2019-10-01T03:23:17-0400

Solution:

Mass percent concentration at 25% means that for each 100g of solution, mass of solute is 25g and mass of solvent is 75 g (1:3 ratio)

"Molality=\frac{moles\,of\, solvent}{1\,kg\,of\,solute}"

Mass of glucose that corresponds to 1 kg of solute is "\frac{1\,kg}{3}=333.3g"

Moles of glucose: "\frac{333.3g}{180.156 g/mol}=1.85 mol"

Molality=1.85 mol/kg

Moles of 1kg of water: "\frac{1000g}{18.015 g/mol}=55.51 mol"

"Mole\,fraction=\frac{moles\,of\,solvent}{moles\,of\,solvent+moles\,of\,solute}"

"Mole\,fraction=\frac{1.85\,mol}{1.85\,mol+55.51\,mol}=0.032"

Mole fraction=0.032